$SPAD/src/input tuplebug.input

Timothy Daly December 3, 2016

Abstract

Contents

2

— tuplebug.spad —

)abbrev package BUG Bug
Bug(R:Ring): with
  bug: R -> (Tuple(R) -> Tuple(R))
 == add
  bug(dummy:R): Tuple(R) -> Tuple(R) == p +-> p

———-

— * —

)set break resume
)spool tuplebug.output
)set message test on
)set message auto off
)clear all
)sys cp $AXIOM/../../src/input/tuplebug.input.pamphlet .
)lisp (tangle "tuplebug.input.pamphlet" "tuplebug.spad" "tuplebug.spad")
--S 1 of 1
)co tuplebug
--R
--R   Compiling AXIOM source code from file
--I      /research/test/int/input/tuplebug.spad using old system compiler.
--R   BUG abbreviates package Bug
--R------------------------------------------------------------------------
--R   initializing nrlib BUG for Bug
--R   compiling into nrlib BUG
--R   compiling exported bug : R -> Tuple R -> Tuple R
--ITime: 0.01 SEC.
--R
--I(time taken in buildFunctor:  0 . NIL)
--R
--R;;;     ***       |Bug| REDEFINED
--R
--R;;;     ***       |Bug| REDEFINED
--ITime: 0 SEC.
--R
--R
--R   Warnings:
--R      [1] bug:  p has no value
--R
--R
--R   Cumulative Statistics for Constructor Bug
--I      Time: 0.01 seconds
--R
--R   finalizing nrlib BUG
--R   Processing Bug for Browser database:
--R--->-->Bug((bug ((Mapping (Tuple R) (Tuple R)) R))): Not documented!!!!

--R--->-->Bug(constructor): Not documented!!!!
--R--->-->Bug(): Missing Description
--R------------------------------------------------------------------------
--I   Bug is now explicitly exposed in frame frame0
--R   Bug will be automatically loaded when needed from
--I      /research/test/int/input/BUG.nrlib/code
--R
--E 1
)spool
)lisp (bye)

———-

References^{[1] nothing}

5